OpenStudy (anonymous):

What is the area of a circle with diameter 8 inches? 4piin2 8piin2 16piin2 64piin2 @bradely

OpenStudy (anonymous):

@dg2

OpenStudy (paki):

use formula... A = π r²

OpenStudy (paki):

and here radius (r)= d/2.... now can you solve...?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Wait yes.

OpenStudy (paki):

hmmm.. show your work here please... ?

OpenStudy (anonymous):

r=4?

OpenStudy (anonymous):

Diameter is 8 correct?

OpenStudy (paki):

r = 8 inches/2 = 4 inches

OpenStudy (paki):

yes,,, then,,,.?

OpenStudy (anonymous):

So it would be pi multiplied by 4?

OpenStudy (anonymous):

3.14 x 4?

OpenStudy (anonymous):

Or 4^2?

OpenStudy (paki):

A = π r² A= 3.14 * (4)^2..... rite... what you got now....

OpenStudy (anonymous):

50.24?

OpenStudy (paki):

oh pie is over there... so remove pie value....

OpenStudy (paki):

A= pie * (4)^2.

OpenStudy (paki):

in the answer options there is a pie, so we will keep as it as.... ok ...

OpenStudy (paki):

A= pie * (4)^2....... solve this... what you got...?

OpenStudy (anonymous):

4^2 = 16 right?

OpenStudy (paki):

yeah good...

OpenStudy (anonymous):

So pie = 3.14?

OpenStudy (paki):

A= pie * (4)^2. A= 16 π inch^2 is the answer....

OpenStudy (anonymous):

r=4 so pi.*16

OpenStudy (anonymous):

area=3.14*4*4=

OpenStudy (anonymous):

50.24

OpenStudy (paki):

no....

OpenStudy (anonymous):

nope.put pi*16

OpenStudy (anonymous):

50.72 mb.

OpenStudy (paki):

@Freelance1 follow, what i have done...

OpenStudy (anonymous):

Is it 50.72?

OpenStudy (paki):

A= pie * (4)^2. A= 16 π inch^2

OpenStudy (anonymous):

third one 16pi in 2

OpenStudy (paki):

what...?

OpenStudy (anonymous):

You both are saying different things.

OpenStudy (anonymous):

third one is the answer

OpenStudy (paki):

we are same....

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

16piin2

OpenStudy (paki):

i have also the same answer.. .check....

OpenStudy (anonymous):

@paki is correct

OpenStudy (anonymous):

Wait mb mb I didn't get what you were trying to say cause I was looking at the wrong question.

OpenStudy (paki):

hahahahah :)

OpenStudy (anonymous):

Can you help me with 3 more?

OpenStudy (paki):

sure... close this one now....